The perimeter of a swimming pool is 91m. The width is 5m more than half the length of the pool. Find the length and width of the pool.
Answer by mananth(16949)

let length be x m
width = x/2 +5 m
...
Perimeter = 2(l+w)
2*(x+x/2+5)= 91
2x+x+10=91
3x+10=91
-10
3x=91-10
3x=81
/3
x=27 m ... the length
...
width = x/2 +5
=27/2 + 5
=13.5 +5
=18.5 m the width
...
CHECK
2*(18.5+27)=91
...
m.ananth@hotmail.ca